OpenStudy (anonymous):

Use the function rule to complete the table. –10x + y = 4

OpenStudy (anonymous):

x= -2, -1, 0, 1, 2 y=

OpenStudy (anonymous):

@MATTW20 can u help me with this one

OpenStudy (anonymous):

@ajprincess can you help mee with this one

OpenStudy (ajprincess):

if u put -2 in place of x in –10x + y = 4 and solve for y what do u get?

OpenStudy (anonymous):

umm not sure

OpenStudy (anonymous):

so what numbers do u have to put in the blank

OpenStudy (anonymous):

wait so does that mean I have to do like -2 + 4 then -1 +4

OpenStudy (ajprincess):

u have to find the values of y. for that plug in each x values given in the equation -10x+y=4 and solve for y. -10x+y=4 y=10x+4 The first value given for x is -2. let us plug in that value in y=10x+4 and find y. y=10(-2)+4 =-20+4 =-16 similarly plug in the other values and find y. Does that help?

OpenStudy (anonymous):

a lil bit ,so the first 2 numbers in y is -2 and -16?

OpenStudy (ajprincess):

only -16

OpenStudy (anonymous):

what u mean only -16 I thought u said the first number is -2

OpenStudy (anonymous):

opps o u said that's for x srry

OpenStudy (ajprincess):

that's k:)

OpenStudy (anonymous):

so how do u find the other 4 numbers

OpenStudy (ajprincess):

what is the second value given for x?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so the next value given for y would be -15?

OpenStudy (ajprincess):

ya. nw plug in that value in y=10x+4 =10(-1)+4 =-10+4

OpenStudy (anonymous):

-6

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

so -16 then -6

OpenStudy (ajprincess):

ya.

OpenStudy (anonymous):

then whats the next 3 numbers

OpenStudy (ajprincess):

plug in the next given value of x in y=10x+4 and find y

OpenStudy (anonymous):

the next value in x is 0

OpenStudy (ajprincess):

ya. what do u get when u plug it in y=10x+4 and solve

OpenStudy (anonymous):

so do 10 x 0 + 4

OpenStudy (ajprincess):

ya

OpenStudy (anonymous):

ok -4

OpenStudy (ajprincess):

check again dr

OpenStudy (anonymous):

4

OpenStudy (ajprincess):

ya good:)

OpenStudy (anonymous):

so now do 10 x 1 +4

OpenStudy (ajprincess):

yup

OpenStudy (anonymous):

14

OpenStudy (anonymous):

now do 10 x 2 +4

OpenStudy (anonymous):

24

OpenStudy (ajprincess):

ya very good:)

OpenStudy (anonymous):

ty I got it now can I ask u another question

OpenStudy (ajprincess):

ya:)

OpenStudy (anonymous):

ok ty .... what is the function rule? x= -2, -1,0,1,2 y= 9, 4, -1, -6, -11

OpenStudy (ajprincess):

let us assume that the function rule is y=mx+c. then plug in 0 for x and -1 for y and find c

OpenStudy (anonymous):

huh? I don't get it

OpenStudy (ajprincess):

let y=mx+c be the equation. (jst an assumption) We need to find m and c. following it?

OpenStudy (anonymous):

i thought we only have to find the function rule

OpenStudy (anonymous):

may u plz just tell me the answer before my test shut off

OpenStudy (ajprincess):

well r u doing an online test?

OpenStudy (anonymous):

yea but not really a test its a 7 question assesment

OpenStudy (anonymous):

and i been on it for an hour and if u take really long when u go press finish it will tell u to log back in and start over

OpenStudy (ajprincess):

well I am sorry. I cannot give you direct answers. It is against the coc of open study

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so how do u find the answers

OpenStudy (anonymous):

@godgavemeyou do u think u can help us with this question

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

what is the function rule? x= -2, -1,0,1,2 y= 9, 4, -1, -6, -11

OpenStudy (anonymous):

whats unit and lesson

OpenStudy (anonymous):

unit 3 lesson 4 science7b

OpenStudy (anonymous):

sciencce?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

srry im just on that quiz

OpenStudy (anonymous):

opss srry im on math

OpenStudy (anonymous):

sry i meant math

OpenStudy (anonymous):

lol im doing the test right now...Imsorry

OpenStudy (anonymous):

im stuck on qestion two lol

OpenStudy (anonymous):

aww man i did all the questions but the last one

OpenStudy (anonymous):

can u give me the answer to 6 and 4

OpenStudy (anonymous):

2 is a

OpenStudy (aravindg):

NO CHEATING IN TEST ALLOWED

OpenStudy (aravindg):

we will alert your teacher if you continue doing the same

OpenStudy (anonymous):

sorry i just need help on this one

OpenStudy (anonymous):

please!!!

OpenStudy (aravindg):

@Callisto , @amistre64

OpenStudy (aravindg):

sorry we take cheating seriously ...

OpenStudy (anonymous):

what ever there happy

OpenStudy (aravindg):

I will help you with every question :) ....after an hour

OpenStudy (aravindg):

let the test get over for now